OpenStudy (anonymous):

Find tan (A - B) for sin A = 12/13 with 0 < A < π/2 and cos B = 8/17 with 0 < B < π/2

OpenStudy (anonymous):

Hello! For some reason, my last question was closed unexpectedly. I may have done it but I do not recall. If I did, it was a total accident. I would appreciate any help on these review questions: http://imgur.com/eBAAHd2 We're studying for the upcoming final and I'm super nervous! Help would be much appreciated as I could really use some!

OpenStudy (anonymous):

Here is some of my work. I just want to know if it's correct before I move on any further: So, sin(A)=12/13 is the same as arcsin(12/13)=A. The arcsin of (12/13) is about equal to (~=) 1.18 radians. Now do the same thing for B: cos(B)=8/17 arccos(8/17)=B 1.08~=B So now we have: tan(1.18-1.08) tan(0.1) ~=0.1